At the end of September, a barrel of light crude oil sold for almost $70 compared to a price near $30 a barrel in January 9f 2004. To answer the following questions,assume bind traders expect inflation to rise from 3% in 2005 to 5% in both 2006 and 2007. Also traders expect the American economy to enter a recession in 2007. Assume the prior to the recent run up oil prices, bond traders had expected inflation to remain stable in 2006 and 2007 at 3%.

Write down the equation representing the Liquidity Premium theory of the terms structure of interest rate. Based on this theory, explain how the yields on short term and medium term government bonds are related,based on the scenerial above.
1
Expert's answer
2021-09-12T19:29:44-0400

Liquidity Premium= the average of past rates and subtract the current rate from that average .

"=\frac{5+3}{2}=4\\4-3=1"


Bonds and interest rates have an inverse relationship. When the cost of borrowing money rises due to high interest rates , bond prices usually fall.The higher a bond's price, the lower its yield